### v1

# Matryoshka: Non-Exclusive Memory Tiering via Transactional Page Migration

Abstract
With the advent of byte-addressable memory devices, such as CXL memory, persistent memory, and storage-class memory, tiered memory systems have become a reality. Page migration is the de facto method within operating systems for managing tiered memory. It aims to bring hot data whenever possible into fast memory to optimize the performance of data accesses while using slow memory to accommodate data spilled from fast memory. While the existing research has demonstrated the effectiveness of various optimizations on page migration, it falls short of addressing a fundamental question: Is exclusive memory tiering, in which a page is either present in fast memory or slow memory, but not both simultaneously, the optimal strategy for tiered memory management?
We demonstrate that page migration-based exclusive memory tiering suffers significant performance degradation when fast memory is under pressure. In this paper, we propose non-exclusive memory tiering, a page management strategy that retains a copy of pages recently promoted from slow memory to fast memory to mitigate memory thrashing. To enable non-exclusive memory tiering, we develop Matryoshka, a new page management mechanism for Linux that features transactional page migration and page shadowing. Matryoshka helps remove page migration off the critical path of program execution and makes migration completely asynchronous. Evaluations with carefully crafted micro-benchmarks and real-world applications show that Matryoshka is able to achieve up to 6x performance improvement over the state-of-the-art transparent page placement (TPP) approach in Linux when under memory pressure. We also compare Matryoshka with a recently proposed access sampling-based page migration approach and demonstrate Matryoshka’s strengths and potential weaknesses in various scenarios. Through the evaluation of Matryoshka, we discover a serious issue facing all tested approaches, unfortunately including Matryoshka, and call for further research on tiered memory-aware memory allocation.
1 Introduction
As new memory devices, such as high bandwidth memory (HBM) [4, 27], DRAM, persistent memory [7, 36], Compute Express Link (CXL)-based memory [1, 41, 34], and storage-class memory [52, 50] continue to emerge, future computer systems are anticipated to feature multiple tiers of memory with distinct characteristics, such as speed, size, power and cost. Tiered memory management aims to leverage the strength of each memory tier to optimize the overall data access latency and bandwidth. Central to tiered memory management is page management within operating systems (OS), including page allocation, placement, and migration. Efficient page management in the OS is crucial for optimizing memory utilization and performance while maintaining transparency for user applications.
Traditionally, the memory hierarchy consists of storage media with at least one order of magnitude difference in performance. For example, in the two-level memory hierarchy assumed by commercial operating systems for decades, DRAM and disks differ in latency, bandwidth, and capacity by 2-3 orders of magnitude. Therefore, the sole goal of page management is to keep hot pages in, and maximize the hit rate of the “performance” tier (DRAM) and migrate (evict) cold pages to the “capacity” tier (disk) when needed. As new memory devices emerge, the performance gap in the memory hierarchy narrows. Evaluations on Intel’s Optane persistent memory  [53] and CXL memory [47] reveal that these new memory technologies are able to achieve comparable performance to DRAM in both latency and bandwidth, within a range of 2-3x. As a result, the assumption on the performance gap, which has guided the design of OS page management for decades, may not hold. It is no longer beneficial to promote a hot page to the performance tier if the migration cost is too high.
Furthermore, unlike disks which must be accessed through the file system as a block device, new memory devices are byte-addressable and can be directly accessed by the processor via ordinary load and store instructions. Therefore, for a warm page on the capacity tier, accessing the page directly and avoiding migration could be a better choice. Most importantly, while the performance of tiered memory remains hierarchical, the hardware is no longer hierarchical. Both the Optane persistent memory and CXL memory appear to the processor as CPUless memory nodes, enabling the OS to manage them as ordinary DRAM nodes.
These unique challenges facing emerging tiered memory systems have inspired research on improving page management in the OS. Much focus has been on expediting page migrations between memory tiers. Nimble [51] improves page migration by utilizing transparent huge pages (THP), multi-threaded migration of a page, and concurrent migration of multiple pages. Transparent page placement (TPP) [41] extends the existing NUMA balancing scheme in Linux to support asynchronous page demotion and synchronous page promotion between fast and slow memory. Memtis [34] and TMTS [22] use hardware performance counters to mitigate the overhead of page access tracking and use background threads to periodically and asynchronously promote pages.
However, these approaches have two fundamental limitations. First, the existing page management for tiered memory assumes that memory tiers are exclusive to each other – hot pages are allocated or migrated to the performance tier while cold pages are demoted to the capacity tier. Therefore, each page is only present in one tier. As memory tiering seeks to explore the tradeoff between performance and capacity, the working set size of workloads that benefit most from tiered memory systems likely exceeds the capacity of the performance tier. Exclusive memory tiering inevitably leads to excessive hot-cold page swapping or memory thrashing when the performance tier is not large enough to hold hot data.
Second, there is a lack of an efficient page migration mechanism to support tiered memory management. As future memory tiers are expected to be addressable by the CPU, page migrations are similar to serving minor page faults and involve three steps: 1) ummap a page from the page table; 2) copy page content to a different tier; 3) remap the page on the page table, pointing to the new memory address. Regardless whether page migration is done synchronously upon accessing a hot page in the slower capacity tier or asynchronously in the background, the 3-step migration process is expensive. During the migration, an unmapped page cannot be accessed by user programs. If page migration is done frequently, e.g., due to memory thrashing, user-perceived bandwidth, including accesses to the migrating pages, is significantly lower (up to 95% lower) than the peak memory bandwidth [51].
This paper advocates non-exclusive memory tiering that allows a subset of pages on the performance tier to have shadow copies on the capacity tier  [^1] We assume that page migrations only occur between two adjacent tiers if there are more than two memory tiers.. Note that non-exclusive tiering is different from inclusive tiering which strictly uses the performance tier as a cache of the capacity tier. The most important benefit is that under memory pressure, page demotion is made less expensive by simply remapping a page if it is not dirty and its shadow copy exists on the capacity tier. This allows for smooth performance transition when memory demand exceeds the capacity of the performance tier.
To reduce the cost of page migration, especially for promotion, this paper proposes transactional page migration (TPM), a novel mechanism to enable page access during migration. Unlike current page migrations, TPM starts page content copy without unmapping the page from the capacity tier so that the migrating page is still accessible by user programs. After page content is copied to a new page on the performance tier, TPM checks whether the page has been dirtied during the migration. If so, the page migration (i.e., the transaction) is invalidated and the copied page is discarded. Failed page migrations will be retried at a later time. If successful, the copied new page is mapped in the page table and the old page is unmapped, becoming a shadow copy of the new page.
We have developed Matryoshka, a new page management mechanism for tired memory that integrates non-exclusive memory tiering and transactional page migration. Matryoshka safeguards page allocation to prevent out-of-memory (OOM) errors due to page shadowing. When the capacity tier is under memory pressure, Matryoshka prioritizes the reclamation of shadow pages before evicting ordinary pages. We have implemented a prototype of Matryoshka in Linux and performed a thorough evaluation. Experimental results show that, compared to two representative page management schemes: TPP and Memtis, Matryoshka achieves up to 6x performance improvement over TPP during memory thrashing and consistently outperforms Memtis by as much as 130% when the working set size fits into fast memory. However, we also discover a potential problem that prevents applications from expanding memory allocations beyond fast memory. This finding motivates more research on the memory allocator in the OS to truly embrace tiered memory.
2 Motivation and Related Work
We introduce the background of page management in tiered memory systems and use TPP [41], a state-of-the-art page placement system designed for CXL-enabled tiered memory, as a motivating example to highlight the main limitations in current page management approaches.
Figure 1: The comparison of memory bandwidth w/ and w/o TPP upon a CXL-based tiered memory: 16 GB DRAM and 16 GB CXL memory. The workload keeps writing memory by following a Zipfian distribution with two working set sizes (WSS): 10 GB and 24 GB. For the 10GB WSS: 6GB of data on DRAM and 4GB on CXL memory, with 10GB of DRAM pre-allocated to other processes (not involved during testing). For the 24GB WSS: 16GB of data on DRAM and 8GB on CXL memory. TPP in progress shows the performance when TPP starts optimizing memory placement (i.e., during the first round of memory accesses); TPP stable displays the performance once TPP has stabilized (after 4 rounds). High spatial locality: Most frequently accessed data is stored on DRAM. Low spatial locality: Frequently accessed data is evenly distributed between DRAM and CXL memory.
Figure 2: Time breakdown during the initial phase of TPP (i.e., TPP in progress): A significant portion of runtime ($\sim$ 34%) is consumed by TPP’s synchronous page migration.
2.1 The Memory Hierarchy
Caching and tiering are two traditional software approaches to manage a memory, or storage hierarchy, consisting of various types of storage media (e.g., CPU caches, DRAM, and hard disks) differing in performance, capacity, and cost. Without loss of generality, we consider a two-layer memory hierarchy with 1) a performance layer (i.e., fast tier), backed with smaller, faster, but more expensive storage media; and 2) a capacity layer (i.e., slow tier) with larger, slower, and cheaper storage media. For caching, data is stored in the capacity layer, and copies of frequently accessed or “hot” data are strategically replicated to the performance layer. For tiering, frequently accessed data is allocated to the performance layer, while less accessed data is relegated to the capacity layer. At any moment, a unique data copy is maintained exclusively.
Caching and tiering share the same goal – to ensure most accesses occur at the performance layer. But, they differ in how they manage data, in the form of pages within an OS. Caching relies on page replacement algorithms to detect hot (or cold) pages and accordingly populate (or evict from) the performance layer while tiering migrates pages between the performance/capacity layers based on the observed page access pattern (hot or cold). Essentially, caching operates in an inclusive page placement mode and retains pages in its original location, merely storing a copy in the performance layer temporarily for fast access. Conversely, tiering operates in an exclusive mode, actively relocating pages across various memory/storage mediums.
Diverse memory/storage devices, such as high bandwidth memory (HBM) [4], CXL-based memory [1], persistent memory (PM) [7], and fast, byte-addressable NVMe SSDs [28], have emerged recently. While they still make a tradeoff between speed, size, and cost, the distinction between them narrows. For example, Intel Optane DC persistent memory (PM), available in a DIMM package on the memory bus enabling programs to directly access data in non-volatile memory from the CPU using load and store instructions, provides an order of magnitude higher capacity than DRAM (e.g., 8x) and offers performance within an order of magnitude of DRAM, e.g., write latency as low as 80 ns and read latency around 170 ns [53]. More recently, CXL, an open, industry-supported interconnect based on PCI Express (PCIe) interface [1], offers a general interface for disaggregating various types of storage devices (e.g., DRAM, PM, and PCIe devices) to the CPU with memory-like, byte-addressable interface (i.e., via the CXL.mem protocol); it allows cache-line granularity access to the connected devices and maintains data coherency and consistency by the underlying hardware. Real-world CXL-based memory offers comparable memory access latency (<2x) and throughput ($\sim$50%) to local DRAM [47].
CXL memory (or PM), with high throughput and low latency, is often treated as an extension of local DRAM, similar to multi-socket Non-Uniform Memory Access (NUMA) nodes. Hence, a tiering based approach is commonly used to manage such a memory hierarchy. Indeed, state-of-the-art tiered memory systems, such as TPP [41], Memtis [34], Nimble [51], and AutoTiering[29], unanimously adopt tiering to exclusively manage data on different memory tiers, i.e., via dynamically migrating pages between fast/slow tiers to ensure efficient access to faster local DRAM.
However, CXL-based tiered memory has its unique characteristics. CXL memory dramatically increases the capacity of local DRAM, accommodating a broader array of applications with data distributed across both fast and slow memory tiers. As data temperature changes over time [41], more frequent data migrations would occur in CXL-based memory tiers. It is especially true when the fast-tier memory is under pressure and incapable of holding all hot data. A high frequency of data migration can cause memory thrashing, hurting overall system performance. As depicted in Figure 1, the involvement of TPP [41] in page migration (i.e., TPP in progress) has a significant, negative impact on the performance of a memory-intensive workload. In comparison, when most of the hot data have been moved to the fast tier (i.e., TPP stable with the 10GB WSS), the workload achieves higher performance than the case without TPP (i.e., No migration). However, if the hot data is larger than the size of the fast tier (the 24GB WSS case), memory thrashing happens, leading to extremely poor performance (i.e., TPP stable with the 24GB WSS). Furthermore, given that CXL-based lower-tier memory can substantially exceed the size of local DRAM [47], the advantage of storing data exclusively for the sake of efficient memory usage becomes less compelling. Instead, more focus should be placed on reducing the overhead when frequent page migration occurs.
Insight 1: In tiered memory systems, adopting an exclusive, NUMA-like approach for data management can be less effective, particularly when the fast-tier memory cannot accommodate all the hot data.
2.2 Page Management
Page management has been widely studied to optimize data placement among tiered memory/storage. An effective page management scheme must tackle two primary challenges: 1) transparent and accurate identification of the hotness of pages with minimal overhead used for page placement; and 2) efficient migration of pages between fast and slow tiers without adversely impacting application-level performance.
Tracking memory access can be conducted by software (via the kernel) and/or with hardware assistance. Specifically, the kernel can keep track of page accesses via page faults [2, 41, 29], scanning page tables [17, 2, 40, 51, 14], or both. Capturing each memory access for precise tracking can be prohibitively slow. For example, the page-fault-based tracking approach traps each memory access to selected pages (i.e., whose page table entry permissions are set to no access) via hint page fault. Thus, it allows the kernel to obtain the accurate recency and frequency of each page. However, page faults reside on the critical path of memory accesses, incurring high latency. On the other hand, the scanning-based approach periodically scans page table entries (PTE) and relies on the reference bit to determine which page has been accessed between scanning intervals. However, it monitors page accesses at a coarser granularity, i.e., a longer scanning interval fails to differentiate the frequency and recency of pages, while a short interval incurs high CPU overhead (up to 70% [34]).
To reduce such overhead, Linux adopts a lazy scanning mechanism, which defers the scanning until memory is tight, e.g., when the fast-tier memory is under pressure, and the swapping mechanism is invoked for page demotion (e.g., via kswapd). In addition, the kernel maintains two separate LRU lists for each memory tier: an active list to store hot pages and an inactive list to store cold pages. Pages can be moved between these two lists during (lazy) scanning. The kernel uses two bits (i.e., PG_referenced and PG_active) to maintain the frequency of each page – a cold page, after two accesses (detected by two scans), will be moved to the active list. Due to lazy scanning, it takes a long time to detect hot pages – i.e., lack of recency. However, timely detecting hot pages for the slow tier is critical. TPP [41] tackles this problem by tracking page accesses of the slow tier using hint page faults. However, we observed that TPP’s synchronous page migration approach unexpectedly amplifies its page-fault overhead, e.g., up to 15 page faults were involved in migrating a single detected hot page from the slow to the fast tiers. This is manifested in Figure 2 – a significant portion of time is consumed by “non-migration page faults” (more details in 3.1).
Accurate and lightweight memory access tracking can be achieved with hardware support, e.g., by adding a PTE count field that records the number of memory accesses (by hardware) for the associated page [42]. However, the proposed approach could be too invasive for mainstream architectures (e.g., x86). In practice, the hardware-assisted sampling approach, via Processor Event-Based Sampling (PEBS) [34, 22], has been recently proposed to obtain page access (virtual address) information from sampled hardware events (e.g., LLC misses or store instructions). However, PEBS-based profiling also requires a careful balance between the frequency of sampling and the accuracy of profiling. We observed that the PEBS-based approach [34], with a sampling rate incurring low CPU overhead, remains coarse-grained and fails to capture many hot pages. Further, the sampling-based approach may not obtain recency of memory access (with a low sampling rate), thus limiting the ability to make timely migration decisions. Last, PEBS is not accessible across CPU vendors [22], making the PEBS-based profiling not generic.
Insight 2: A practical, lightweight memory tracking mechanism that can capture both frequency and recency information of page accesses is highly desired.
Page migration – i.e., the process of promoting hot pages from the slow tier to the fast tier or demoting cold pages from the fast tier to the slow tier – involves a complex procedure: ① The system needs to trap to the kernel (e.g., via page faults); ② The PTEs of selected pages must be locked (to avoid others accessing the pages during migration) and ummapped; ③ A translation lookaside buffer (TLB) shootdown must be issued to each processor (via IPI) that may have cached copies of stale PTEs; ④ The content of the pages are copied between tiers; ⑤ Finally, their corresponding PTEs must be updated and remapped. In addition, page migration can be initiated in a synchronously or asynchronously manner. The triggering mode is also linked with the memory tracking mechanisms.
Typically, scanning-based tracking supports asynchronous, batch-based page promotion/demotion, which redistributes pages among fast/slow tiers based on the periodic updates of data temperature [34]. While batching multiple pages for asynchronous migration could reduce migration overhead, it does not favor temporal locality. To migrate on-demand pages more promptly, page-fault-based, synchronous page migration is often used in both Linux and recent research projects [41, 51]: When a “tagged” page is touched for the first time, page migration is synchronously conducted during handling the hint page fault. Despite being timely, synchronous migration is very costly. First, it resides in the critical path of memory access, slowing down the execution of the process requesting such memory access. Further, due to locking (in ①), others trying to access the same page must enter a busy-wait state until the page becomes available remapped). Last, page migration may fail, e.g., due to insufficient space on the destination memory tier, and the need to synchronously re-try further worsens this situation.
This exactly explains the extremely poor performance caused by TPP, as shown in Figure 1, i.e., the cases in “TPP in progress” and “TPP stable” with the 24 GB WSS. The breakdown of time consumption, in Figure 2, further shows that almost 30% of the execution time was spent in the synchronous page promotion (and related) operations, significantly slowing down the workload.
Insight 3: How can we achieve prompt page migration without delaying/pausing the execution of applications?
2.3 Related Work
A long line of pioneering work has explored a wide range of of tiered storage/memory systems, built upon SSDs and HDDs [12, 15, 20, 54, 46, 38, 49, 30, 24], DRAM and disks [19, 23, 26, 43], HBM and DRAMs [42, 21, 45], NUMA memory [2, 3], PM and DRAM [37, 13, 40, 14], local and far memory [17, 25, 31, 32, 44], DRAM and CXL memory [35, 41, 34], and multiple tiers [48, 29, 51, 37, 33]. We focus on tiered memory systems consisting of DRAM and emerging CXL memory, though the proposed approaches in Matryoshka also apply to other tiered memory systems such as HBM/DRAM and DRAM/PM. We discuss the most related works as follows.
Lightweight memory access tracking: To mitigate software overhead associated with memory access tracking, Hotbox [17] employs two separate scanners for fast and slow tiers to scan the slow tier at a fixed rate while the local tier at an adaptive rate, configurable based on the local memory pressure. Memtis [34] adjusts its PEBS-based sampling rate to ensure its overhead is under a threshold (e.g., < 3%). TMTS [22] also adopts a periodic scanning mechanism to detect frequency along with hardware sampling to more timely detect newly hot pages. While these approaches balance scanning/sampling overhead and tracking accuracy, an “always-on” profiling component does not seem practical, especially for high-pressure workloads. Instead, thermostat [14] samples a small fraction of pages, while DAMON [3] monitors memory access at a coarser-grained granularity (i.e., region). Although both can effectively reduce the scanning overhead, coarse granularity leads to lower accuracy regarding page access patterns. On the other hand, to reduce the overhead associated with frequent hint page faults like AutoNUMA [2], TPP [41] enables the page-fault based detection only for CXL memory (i.e., the slow tier) and tries to promote a page promptly (e.g., when it encounters a page fault); prompt page promotion avoids subsequent page faults on the same page. Inspired by existing lightweight tracking systems, such as Linux’s active and inactive lists and hint page faults, Matryoshka advances them by incorporating more recency information with no additional CPU overhead. Unlike hardware-assisted approaches [34, 22, 39], Matryoshka does not require any additional hardware support.
Page migration optimizations: To hide reclamation overhead from applications, TPP [41] decouples page allocation and reclamation; however, page migration remains in the critical path, incurring significant slowdowns. Nimble [51] focuses on mitigating page migration overhead with new migration mechanisms, including transparent huge page migration and concurrent multi-page migration. Memtis [34] further moves page migration out of the critical path using a kernel thread to promote/demote pages in the background. TMTS [22] leverage a user/kernel collaborative approach to control page migration. In contrast, Matryoshka aims to achieve prompt, on-demand page migration while moving page migration off the critical path. It is orthogonal to and can benefit from existing page migration optimizations. The most related work is [18], which leverages hardware support to pin data in caches, enabling access to pages during migration. Again, Matryoshka does not need additional hardware support.
3 Matryoshka Design and Implementation
Matryoshka is a new page management mechanism for tiered memory that features non-exclusive memory tiering and transactional page migration. The goal of Matryoshka design is to enable the processor to freely and directly access pages from both fast and slow memory tiers and move the cost of page migration off the critical path of users’ data access. Note that Matryoshka does not make page migration decisions and relies on memory tracing in the OS to determine hot and cold pages. Furthermore, Matryoshka does not impact the initial memory allocation in the OS and assumes a standard page placement policy. Pages are allocated from the fast tier whenever possible and are placed in the slower tier only when there is an insufficient number of free pages in the fast tier, and attempts to reclaim memory in the fast tier have failed. After the initial page placement, Matryoshka gradually migrates hot pages to the fast tier and old pages to the slow tier. Matryoshka seeks to address two key issues: 1) how to minimize the cost of page migration? 2) how to minimize the number of migrations?
Overview. Inspired by multi-level cache management in modern processors, which do not employ a purely inclusive or exclusive caching policy between tiers [16] to facilitate the sharing of or avoid the eviction of certain cache lines, Matryoshka embraces a non-exclusive memory tiering policy to prevent memory thrashing when under memory pressure. Unlike the existing page management schemes that move pages between tiers and require that a page is only present in one tier, Matryoshka instead copies pages from the slow tier to the fast tier and keeps a shadow copy of the migrated pages at the slow tier. The non-exclusive tiering policy maintains shadow copies only for pages that have been promoted to the fast tier, thereby not an inclusive policy. The advantage of the non-exclusive policy is that the demotion of clean, cold pages can be simplified to remapping the page table entry (PTE) without the need to actually copy the cold page to the slower tier.
The building block of Matryoshka is a new transactional page migration (TPM) mechanism to reduce the cost of page migrations. Unlike the existing unmap-copy-remap 3-step page migration, TPM opportunistically copies a page without unmapping it from the page table. During the page copy, the page is not locked and can be accessed by a user program. After the copy is completed, TPM checks if the page has been dirtied during the copy. If not, TPM locks the page and remaps it in the PTE to the faster tier. Otherwise, the migration is aborted and will be tried at a later time. TPM not only minimizes the duration during which a page is inaccessible but also makes page migration asynchronous, thereby removing it from the critical path of users’ data access.
Without loss of generality, we describe Matryoshka design in the context of Linux. We start with transactional page migration and then delve into page shadowing – an essential mechanism that enables non-exclusive memory tiering.
Figure 3: The workflow of transactional page migration. PFN is the page frame number and D is the dirty bit in PTE. The page is only inaccessible by user programs during step 4 when the page is remapped in the page table.
3.1 Transactional Page Migration
The motivation to develop TPM is to make page migration entirely asynchronous and decoupled from users’ access to the page. As discussed in Section 2.2, the current page migration in Linux is synchronous and on the critical path of users’ data access. For example, the default tiered memory management in Linux, TPP, attempts to migrate a page from the slow tier whenever a user program accesses the page. Since the page is in “protected” mode, the access triggers a minor page fault, leading TPP to attempt the migration. The user program is blocked and makes no progress until the minor page fault is handled and the page is remapped to the fast tier, which can be a time-consuming process. Worse, if the migration fails, the OS remains in function migrate_pages and retries the aforementioned migration until it is successful or reaching a maximum of 10 attempts.
TPM decouples page migration from the critical path of user programs by making the migrating page accessible during the migration process. Therefore, users will access the migrating page from the slow tier before the migration is complete. While accessing a hot page from the slow tier may lead to sub-optimal memory performance, it avoids blocking user access due to the migration, thereby leading to superior user-perceived performance. Figure 3 shows the workflow of TPM. Before migration commences, TPM clears the protection bit of the page frame and adds the page to a migration pending queue. Since the page is no longer protected and it is not yet unmapped from the page table, following accesses to the page will not trigger additional page faults.
TPM starts a migration transaction by clearing the dirty bit of the page (step ❶) and checks the dirty bit after the page is copied to the fast tier to determine whether the transaction was successful. After changing the dirty bit in PTE, TPM issues a TLB shootdown to all cores that ever accessed this page (step ❷). This is to ensure that subsequent writes to the page can be recorded on the PTE. After the TLB shootdown is completed, TPM starts copying the page from the slow tier to the fast tier (step ❸). To commit the transaction, TPM checks the dirty bit by loading the entire PTE using atomic instruction get_and_clear (step ❹). Clearing the PTE is equivalent to unmapping the page and thus another TLB shootdown is needed (step ❺). Note that after unmapping the page from PTE, it becomes inaccessible by users. TPM checks whether the page was dirtied during the page copy (step ❻) and either commits the transaction by remapping the page to the fast tier if the page is clean (step ❼) or otherwise aborts the transaction (step ❽). If the migration is aborted, the original PTE is restored and waits for the next time when TPM is rescheduled to retry the migration. The duration in which the page is inaccessible is between ❹ and ❼/ ❽, significantly shorter than that in TPP (possibly multiple attempts between  ❶ and ❼).
Page migration is a complex procedure that involves memory tracing and updates to the page table for page remapping. The state-of-the-art page fault-based migration approaches, e.g., TPP in Linux [41], employs synchronous page migration, a mechanism in the Linux kernel for moving pages between NUMA nodes. In addition to the extended migration time affecting the critical path of user programs, this mechanism causes excessive page faults when integrated with the existing LRU-based memory tracing. TPP makes per-page migration decisions based on whether the page is on the active LRU list. Nevertheless, in Linux, memory tracing adds pages from the inactive to the active LRU list in batches of 15 requests  [^2] The 15 requests could be repeated requests for promoting the same page to the active LRU list, aiming to minimize the queue management overhead. Due to synchronous page migration, TPP may submit multiple requests (up to 15 if the request queue was empty) for a page to be promoted to the active LRU list to initiate the migration process. In the worst case, migrating one page may generate as many as 15 minor page faults.
TPM provides a mechanism to enable asynchronous page migration but requires additional effort to interface with memory tracing in Linux to minimize the number of page faults needed for page migration. As shown in Figure 4, in addition to the inactive and active LRU lists in memory tracing, TPM maintains a separate promotion candidate queue (PCQ) for pages that 1) have been tried for migration but 2) not yet promoted to the active LRU list. Upon each time a minor (hint) page fault occurs and the faulting page is added to PCQ, TPM checks if there are any hot pages in PCQ that have both the active and accessed bits set. These hot pages are then inserted to a migration pending queue, from where they will be tried for asynchronous, transactional migration by a background kernel thread kpromote. Note that TPM does not change how Linux determines the temperature of a page. For example, in Linux, all pages in the active LRU list, which are eligible for migration, have the two memory tracing bits set. However, not all pages with these bits set are in the active list due to LRU list management. TPM bypasses the LRU list management and provides a more efficient method to initiate page migration. If all transactional migrations were successful, TPM guarantees that only one page fault is needed per migration in the presence of LRU list management.
Figure 4: TPM uses a two-queue design to enable asynchronous page migration.
3.2 Page Shadowing
Figure 5: Shadow page management using shadow r/w bit.
To enable non-exclusive memory tiering, Matryoshka introduces a one-way page shadowing mechanism to enable a subset of pages resident in the performance tier to have a shadow copy in the capacity tier. Only pages promoted from the slow tier have shadow copies in the slow tier. Shadow copies are the original pages residing on the slow tier before they are unmapped in the page table and migrated to the fast tier. Shadow pages play a crucial role in minimizing the overhead of page migration during periods of memory pressure. Instead of swapping hot and cold pages between memory tiers, page shadowing enables efficient page demotion through page table remapping. This would eliminate half of the page migration overhead, i.e., page demotion, during memory thrashing.
Indexing shadow pages. Inspired by the indexing of file-based data in the Linux page cache, Matryoshka builds an XArray for indexing shadow pages. An XArray is a radix-tree like, cache-efficient data structure that acts as a key-value store, mapping from the physical address of a fast tier page to the physical address of its shadow copy on the slow tier. Upon successfully completing a page migration, Matryoshka inserts the addresses of both the new and old pages into the XArray. Additionally, it adds a new shadow flag to the struct page of the new page, indicating that shadowing is on for this page.
Shadow page management. The purpose of maintaining shadow pages is to assist page demotion. Fast or efficient page demotion is possible via page remapping if the master page, i.e., the one on the fast tier, is clean and consistent with the shadow copy. Otherwise, the shadow copy should be discarded. To track inconsistency between the master and shadow copies, Matryoshka sets the master page as read-only and a write to the page causes a page fault. To simplify system design and avoid additional cross-tier traffic, Matryoshka discards the shadow page if the master page is written.
However, tracing updates to the master page poses a significant challenge. Page management in Linux relies heavily on the read-write permission of a page to perform various operations on the page, such as copy-on-write (CoW). While setting master pages as read-only effectively captures all writes, it may affect how these master pages are managed in the kernel. To address this issue, Matryoshka introduces a procedure called shadow page fault. It still designates all master pages as read-only but preserves the original read-write permission in an unused software bit on the page’s PTE (as shown in Figure 5). We refer to this software bit as shadow r/w. Upon a write to a master page, a page fault occurs. Unlike an ordinary page fault that handles write violation, the shadow page fault, which is invoked if the page’s shadow flag is set in its struct page, restores the read-write permission of the faulting page according to the shadow r/w bit and discards/frees the shadow page on the slow tier. The write may proceed once the shadow page fault returns and reinstates the page to be writable. For read-only pages, tracing shadow pages does not impose additional overhead; for writable pages, it requires one additional shadow page fault to restore their write permission.
Reclaiming shadow pages. Non-exclusive memory tiering introduces space overhead due to the storage of shadow pages. If shadow pages are not timely reclaimed when the system is under memory pressure, applications may encounter out-of-memory (OOM) errors, which would not occur under exclusive memory tiering. There are two scenarios in which shadow pages should be reclaimed. First, the Linux kernel periodically checks the availability of free pages and if free memory falls below low_water_mark, kernel daemon kswapd is invoked to reclaim memory. Matryoshka instructs kswapd to prioritize the reclamation of shadow pages. Second, upon a memory allocation failure, Matryoshka also tries to free shadow pages. To avoid OOM errors, the number of freed shadow pages should exceed the number of requested pages. However, frequent memory allocation failures could negatively affect system performance. Matryoshka employs a straightforward heuristic to reclaim shadow pages, targeting 10 times the number of requested pages or until all shadow pages are freed. While excessive reclamation may have a negative impact on Matryoshka’s performance, it is crucial to prevent Out-of-Memory (OOM) errors. Experiments in Section 4 demonstrate the robustness of Matryoshka even under extreme circumstances.
3.3 Limitation
Matryoshka relies on two rounds of TLB shootdown to effectively track updates to a migrating page during transactional page migration. When a page is used by multiple processes or mapped by multiple page tables, its migration involves multiple TLB shootdowns, per each mapping, that need to happen simultaneously. The overhead of handling multiple IPIs could outweigh the benefit of asynchronous page copy. Hence, Matryoshka deactivates transactional page migration for multi-mapped pages and resorts to the default synchronous page migration mechanism in Linux.
4 Evaluation
Figure 6: The three memory provisioning schemes used in evaluating tiered memory management. From bottom to top (with respect to fast memory): over-provisioning, approaching capacity, and under-provisioning.
This section presents a thorough evaluation of Matryoshka on its performance, overhead, and robustness. Most importantly, we aim at understanding tiered memory management through the comparison between Matryoshka and the existing representative approaches in three scenarios. We study two types of memory footprints: 1) resident set size (RSS) – the total size of memory occupied by a program; 2) working set size (WSS) – the amount of memory a program actively uses during execution. Since we focus on in-memory computing, WSS is typically smaller than RSS. Figure 6 shows the three scenarios we study to understand the tradeoff between migrating pages to the fast tier and accessing them directly from the slow tier. RSS determines the initial page placement while WSS decides the number of pages that should be migrated to the fast tier.
Platform A Platform B Platform C
(engineering sample)
CPU 4th Gen Xeon Gold 4th Gen Xeon Platinum 2nd Gen Xeon Gold
DRAM 16 GB DDR5 16 GB DDR5 16 GB DDR4
CXL or PM Memory Agilex 7 16 GB DDR4 Agilex 7 16 GB DDR4 Optane 100 256 GB DDR-T
Table 1: The configurations of three testbeds.
Figure 7: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 10 GB and a total of 20 GB of RSS data.
Figure 8: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 13.5 GB and a total of 27 GB RSS.
Workload Type In progress Promotion In progress Demotion Steady Promotion Steady demotion
Small WSS (2.6M|1.5M)/((1.5M|1.48M) (3.3M|654K)/(1.63M|1.49M) (2|0)/(1|2) (0|0)/(0|0)
Medium WSS (6M|6M)(3.7M|3.9M) (6M|5.8M)/(3.7M|3.9M) (2.1M|2.3M)/(119K|3.35M) (2.1M|2M)/(116K|3.36M)
Large WSS (12.5M|12.3M)/(11.6M|11.5M) (13M|13M)/(12.3M|12M) (11.8M|11.6M)/(7.2M|11.3M) (11.9M|11.7M)/(7.2M|11.3M)
Table 2: The number of page promotions/demotions (read|write) during the initial in-process phase and the stable phase for TPP/Matryoshka. Note that, the number for Memtis is too small compared with TPP/Matryoshka, hence not listed.
Figure 9: Micro-benchmark: performance comparisons between TPP, Memtis, and Matryoshka with a WSS of 27 GB and the same 27 GB RSS.
Workload type Success : Aborted
10GB WSS: read 148 : 1
10GB WSS: write 197 : 1
13.5GB WSS: read 151 : 1
13.5GB WSS: write 267 : 1
27GB WSS: read 167 : 1
27GB WSS: write 153 : 1
Table 3: The success rate of transactional migration.
Experimental setup: We compared Matryoshka with two state-of-the-art tired memory systems: TPP [41] and Memtis [34]. We conducted experiments upon three platforms with different configurations in CPU, local DRAM, CXL memory, and persistent memory, as listed in Table 1. Notably, we used a real-world, Agilex-7 FPGA-based CXL memory device [6] and Intel Optane 100 Series Persistent Memory (PM). The CXL memory device has only 16 GB of memory. Hence, we configured local DRAM with a size of 16 GB  [^3] Owing to the limited availability of CXL memory devices, we could not test Matryoshka with a larger CXL memory tier. However, Matryoshka’s non-exclusive design would benefit more from a large-capacity slow tier.. The size of the Intel Optane PM was 256 GB. In this tiered memory, local DRAM serves as the fast memory tier, while CXL or PM serves as the slow memory tier. In addition, each of the three platforms was equipped with a single socket, and we configured the number of CPU cores to 32. We evaluate both TPP and Matryoshka on Linux kernel v5.15, and run Memtis on kernel v5.19. We first used micro-benchmarks to investigate the performance of Matryoshka’s transactional page migration and page shadowing mechanisms. Further, we used real-world applications to assess the overall performance of Matryoshka.
4.1 Micro-benchmarks
We developed a micro-benchmark to precisely assess Matryoshka in a controlled manner. This micro-benchmark involves 1) allocating data to specific segments of the tiered memory, 2) running with various working set sizes (WSS) and resident set sizes (RSS); and 3) generating memory accesses to the WSS data that mimic real-world memory access patterns with a Zipfian distribution. We created three distinct scenarios, which represent the small, medium, and large WSS, for the micro-benchmarking. We conducted all the micro-benchmarking on Platform B – a tiered memory system with 16 GB local DRAM and 16 GB remote CXL memory.
A small WSS: We began with a scenario by running the micro-benchmark with a small WSS of 10 GB and a total RSS of 20 GB. Initially, we filled the first 10 GB of local DRAM with the first half of the RSS data. Subsequently, we allocated 10 GB of WSS data as the second half of the RSS data: 6 GB on the local DRAM and 4 GB on the CXL memory. The micro-benchmark, during its operation, continuously executed memory reads or writes (following a Zipfian distribution) to this 10 GB WSS data, spread across both the local DRAM and CXL memory. The frequently accessed data, or “hot” data, was uniformly distributed along the 10 GB WSS. For both TPP and Matryoshka, accessing data on the CXL memory triggered its migration to the local DRAM, with TPP doing so synchronously and Matryoshka asynchronously. In contrast, Memtis used a background thread to migrate hot data from the CXL memory to the local DRAM. Due to page migration, the 4 GB data on the CXL memory gradually moved to the local DRAM – i.e., a part of the cold RSS data was demoted to the CXL memory. As the total WSS was small (i.e., 10 GB), it was stored in the local DRAM after the micro-benchmark ran for a while, reaching a stable state.
Figure 7 illustrates that in the initial phase, where intensive page migration was conducted (i.e., migration in progress), both Matryoshka and Memtis demonstrated equivalent performance regarding memory bandwidth for reads and writes. During this period, they outperformed TSS significantly, with improvements ranging from 100% to 300%. The performance benefit came from the asynchronous, transactional page migration for Matryoshka (Section 3.1) and asynchronous, background page migration for Memtis, which merely blocked the execution of the micro-benchmark.
Figure 7 further shows the results in the stable phase (i.e., migration stable), where most of the WSS data was migrated from the CXL memory to the local DRAM. Both Matryoshka and TPP achieved equivalent read/write performance because memory accesses were all served by the local DRAM and there were few or no page migrations, as listed in Table 2. Surprisingly, the performance of Memtis was the worst (less than 40% of the other two). In fact, the performance with Memtis did not change much from the initial phase to the stable phase. Possible reasons include 1) Memtis’s PEBS-based sampling was too coarse-grained and failed to capture most hot pages on the CXL memory; 2) Memtis’s background migration thread could not timely migrate pages from the local DRAM to the CXL memory. We only observed a small number of page migrations occurring during both the initial and stable phases (not listed in Table 2). Hence, in the stable phase, many memory accesses were still served by the slow CXL memory, leading to low performance.
A medium WSS: We continued to increase the WSS and the RSS to a larger size, setting both at 13.5 GB and 27 GB separately. Once again, we placed the first half of the RSS data (13.5 GB) at the start of the local DRAM, followed by 2.5 GB of the WSS on the local DRAM, with the remaining 11 GB residing on the CXL memory.
Similar to the small WSS case, Figure 8 demonstrates that during the initial phase, where heavy page migrations were conducted (migration in progress), both Matryoshka and Memtis achieved comparable memory bandwidths for read and write operations. In this period, they significantly outperformed TPP, showing improvements in the range of 2x to 3x. Compared with the small WSS case, in the medium WSS case, more page migrations were conducted in the initial phase – e.g., more than 2x as listed in Table 2. This is because a larger portion of the WSS data was initially allocated in the CXL memory than the small WSS case (11 GB vs. 4 GB).
Figure 8 presents the findings from the stable phase (migration stable). In this phase, Matryoshka and Memtis exhibited nearly the same performance in terms of write operations. However, for read operations, Matryoshka notably excelled, outperforming the other two approaches substantially – for instance, it achieved 4x the performance of TSS and 2x the performance of Memtis. In the medium WSS case, we observed that, during the stable phase, a significant number of page migrations were still triggered, as listed in Table 2. The reason is that although the WSS of 13.5 GB is smaller than the local DRAM capacity of 16 GB, the combined size of the WSS data and the system’s data exceeds the threshold for space reclaiming, triggering the demotion thread to begin migrating pages to the CXL memory. The demoted pages belonging to the WSS were promptly promoted back to the local DRAM shortly, leading to memory thrashing.
Matryoshka can significantly mitigate the performance degradation for read operations (resulting from memory thrashing) due to its non-exclusive memory tiering. Recall that Matryoshka’s page shadowing mechanism enables a subset of pages resident in the local DRAM to have a shadow copy in the CXL (Section 3.2). This mechanism benefits the read operations when thrashing occurs – when a page is selected for demotion and meanwhile Matryoshka retains its copy in the CXL memory, the page is discarded rather than migrated. Unfortunately for writes, as the local version has been modified during demotion, the new version needs to be migrated to the CXL memory. But Matryoshka still managed to deliver the best performance among the three approaches.
A large WSS: In our final test using micro-benchmarks, we scaled up the WSS (and the RSS) to a size of 27 GB. In this scenario, all the data accessed in the micro-benchmark constituted the WSS data. We arranged this data sequentially, starting from the local DRAM and extending to the CXL memory, allocating 16 GB of the WSS on the local DRAM and the remaining 11 GB on the CXL memory. This created an extreme scenario that caused severe memory thrashing, as the total hot data greatly (27 GB) exceeded the capacity of the local DRAM (16 GB).
Figure 9 presents the performance in the initial phase and the the stable phase. Interestingly, both Matryoshka and TPP, tending to promptly migrate detected hot data from the CXL memory, performed worse than Memtis, which only conducted background promotion/demotion. This is because, the severe memory thrashing caused by Matryoshka and TPP, frequently migrated pages between the local and remote memory, greatly degrading the micro-benchamrk’s performance. In contrast, much fewer page migrations occurred with Memtis. However, Matryoshka consistently outperformed TPP by a range of 2x - 5x, again due to its asynchronous, transactional page migration. In addition, Matryoshka’s page shadowing helped achieve higher performance. Regardless, frequent page migrations become counterproductive when the WSS significantly exceeds the capacity of fast memory.
Migration success rate: As discussed in Section 3.1, there exists a possibility that Matryoshka’s transactional page migration might not succeed due to simultaneous modification (dirtying) of the page by a process. Table 3 demonstrates that, across various scenarios tested with micro-benchmarks, Matryoshka maintains a notably high success rate in page migrations.
RSS 23GB 25GB 27GB 29GB
Total shadow page size 3.93GB 2.68GB 2.2GB 0.58GB
Table 4: Robustness evaluation on Platform B with 30.7GB tiered memory.
Robustness: Page shadowing can potentially increase memory usage and in the worst case can cause OOM errors. In this test, we evaluated the effectiveness of Matryoshka’s shadow page reclamation. We measured the total memory usage and the size of shadow memory using a micro-benchmark that sequentially scans a predefined RSS area. Table 4 shows the change of shadow pages as we varied the RSS. The results suggest that Matryoshka was able to effectively reclaim shadow pages as the benchmark approached the capacity of the total memory. We further pushed the RSS (not shown in the table) beyond the total memory size. The system started swapping but encountered no OOM errors. We repeated the tests for a large number of times and no OOM errors were observed.
Summary: Matryoshka enables applications to sustain high performance via asynchronous, transactional page migrations when the WSS is smaller than the performance tier. Additionally, when the WSS is comparable to or exceeds the capacity of the performance tier, Matryoshka’s page shadowing feature notably enhances the efficiency of read operations.
Figure 10: Performance comparisons using Redis and YCSB between TPP and Matryoshka. Despite our best efforts, we could not execute Memtis on Redis successfully.
Figure 11: Performance comparisons with PageRank between non-migration, TPP, Memtis, and Matryoshka.
Figure 12: Performance comparisons with Liblinear between non-migration, TPP, Memtis, and Matryoshka.
4.2 Real-world Applications
We continued the evaluation of Matryoshka using three representative real-world applications with unique memory access patterns: Redis [10], PageRank [9], and Liblinear [5]. We ran these three applications on top of three different platforms (Table 1) to reveal the benefits and potential limitations of current page management systems for emerging tiered memory.
Key-value store: We first conducted experiments on a latency-sensitive key-value database, Redis [10]. The workload was generated from YCSB [11], using its update heavy workload A, with a 50/50 distribution of read and write operations. We carefully crafted three cases with different RSS and total operations. Note that the parameters of YCSB were set as default if they were not specified as follows.
Case 1: We set recordcount to 6 million and operationcount to 8 million. After pre-loading the dataset into the database, we used a customized tool to demote all Redis’ memory pages to the slow memory tier. Then, we started to run the experiment. The RSS of this case was 13GB.
Case 2: We increased the RSS by setting recordcount to 10 million and operationcount to 12 million. We used the same way to demote all the memory pages to the slow tier. The RSS of this case was 24GB.
Case 3: We increased the total operations while keeping the same RSS by setting recordcount to 10 million and operationcount to 12 million. After pre-loading the dataset, we did not demote any memory pages. The RSS was 24GB.
Figure 10 shows that, compared to TPP, Matryoshka consistently delivered superior performance for the latency-sensitive key-value store across all tested platforms in the three cases. The enhanced performance attributable to Matryoshka stems from its asynchronous, transactional page migration and page shadowing approaches, which merely block the execution of applications. Notably, Matryoshka demonstrated the highest performance in case 3. This is due to that a smaller amount of data (8 GB) was initially allocated on the slower tier in case 3, compared to the larger initial allocations (13 GB and 16 GB) in the other two cases.
Graph-based computation: We further used PageRank [9], an application used to rank web pages. It involves iterative computations to determine the rank of each page, based on the link structures of the entire web. As the size of the dataset increases, the computational complexity also increases, making it both memory-intensive and compute-intensive. We used a benchmark suite [8] to generate a synthetic uniform-random graph comprising $2^{26}$ vertices, each with an average of 20 edges. The RSS in this experiment was 22 GB, indicating that the memory pages were distributed at both the local DRAM and remote CXL memory or PM.
Figure 11 illustrates that there is negligible variance in performance between scenarios with page migrations (using Matryoshka and TPP) and without page migrations (no migration). The results suggest that: 1) For non-latency-sensitive applications, such as PageRank, using CXL memory (as in Platform A) can significantly expand the local DRAM capacity without adversely impacting application-level performance. 2) In such scenarios, page migration appears to be unnecessary. In addition, the application-level performance diminishes with PM (as observed in Platform C), partly due to PM’s higher latency compared to the CXL memory. These findings also reveal that the overhead associated with Matryoshka’s page migration minimally influences PageRank’s performance. Additionally, it is observed that among all evaluated scenarios, Memtis exhibits the least efficient performance.
Machine learning: Our final evaluation of Matryoshka involved using the machine learning library Liblinear [5], known for its large-scale linear classification capabilities. We executed Liblinear with an L1 regularized logistic regression workload, using 16 threads with an RSS of 10 GB. Prior to each execution, we used our tool to demote all memory pages associated with the Liblinear workload to the slower memory tier.
Figure 12 illustrates that both Matryoshka and TPP significantly outperformed the “no migration” scenario on Platform B, which utilizes CXL-based memory tiering. However, this performance advantage was not evident on Platform C, featuring PM-based memory tiering, where all approaches exhibited comparable performance. Further analysis revealed a sharp difference in page promotions between the platforms: 12 million on Platform B compared to just 1 million on Platform C. This discrepancy is attributable to the differing speeds of CXL memory and PM. This observation indicates that the efficacy of page migration is closely linked to the performance characteristics of the underlying memory media.
Our observations from Figure 11 and Figure 12 also reveal that the application-level performance was related to the distinct characteristics of platforms. For instance, Platform A represents a standard off-the-shelf machine, whereas Platform B is an engineering sample with undisclosed features, which may affect various applications differently. This implies that the presence of other prominent performance-related factors could potentially overshadow the effects of page migrations.
5 Discussions and Future Work
The key insight from Matryoshka’s evaluation is that page migration, especially under memory pressure, has a detrimental impact on overall application performance. While Matryoshka achieved graceful performance degradation and much higher performance than TPP, an approach based on synchronous page migration, its performance is sub-optimal compared to that without page migration. When the program’s working set exceeds the capacity of the fast tier, the most effective strategy is to access pages directly from their initial placement, completely disabling page migration. It is straightforward to detect memory thrashing, e.g., frequent and equal number of page demotions and promotions, and disable page migrations. However, estimating the working set size to resume page migration becomes challenging, as the working set now spans multiple tiers. It requires global memory tracking, which could be prohibitively expensive, to identify the hot data set that can potentially be migrated to the fast tier. We plan to extend Matryoshka to unilaterally throttle page promotions and monitor page demotions to effectively manage memory pressure on the fast tier. Note that this would require the development of a new page migration policy, which is orthogonal to the Matryoshka page migration mechanisms proposed in this work.
There exist difficult tradeoffs between page fault-based access tracking, such as TPP and Matryoshka, and hardware performance counter-based memory access sampling like Memtis. While page fault-based tracking effectively captures access recency, it can be potentially expensive and on the critical path of program execution. In comparison, hardware-based access sampling is off the critical path and captures access frequency. However, it is not responsive to workload changes and its accuracy relies critically on the sampling rate. One advantage of Matryoshka is that it is a page fault-based migration approach that is asynchronous and completely off the critical path. A potential future work is to integrate Matryoshka with hardware-based, access frequency tracking, such as Memtis, to enhance the current migration policy.
Lastly, we have observed application and system crashes on platform-C which has almost an order of magnitude more (persistent) memory on the capacity tier than the performance tier. The crashes occurred to all tested approaches including the vanilla Linux kernel, and to applications whose working set size is significantly larger than the capacity of fast memory. Intuitively, the Linux kernel should utilize tiered memory equally when under memory pressure. However, after reviewing system logs and Linux kernel source code, we found that Linux, and provably many other commercial OSes, follows a hierarchical approach in managing the performance and memory allocation of tiered memory – attempting to cache or allocate data from fast memory if at all possible and resorting to slow memory if failed. Although there are guardrails to spill data into slow memory, such as kswapd, these mechanisms are often not responsive enough to prevent an OOM error. It is worth noting that Matryoshka has robust shadow page reclamation and does not cause OOM errors due to page shadowing. This finding motivates further research on tiered memory allocation that takes into account both non-hierarchical strategies for reliability and hierarchical strategies for performance.
6 Conclusion
This paper introduces non-exclusive memory tiering as an alternative to the common exclusive memory tiering strategy, where each page is confined to either fast or slow memory. The proposed approach, implemented in Matryoshka, leverages transactional page migration and page shadowing to enhance page management in Linux. Unlike traditional page migration, Matryoshka ensures asynchronous migration and retains shadow copies of recently promoted pages. Through comprehensive evaluations, Matryoshka demonstrates up to 6x performance improvement over existing methods, addressing critical performance degradation issues in exclusive memory tiering, especially under memory pressure. The paper calls for further research in tiered memory-aware memory allocation.